OpenStudy (anonymous):

pls help me solve xdy/dx-y(+y^2)e^2x=0 using y=v^-1

OpenStudy (anonymous):

\[x \frac{ dy }{ dx}-y+y ^{2}e ^{2x}=0 using y=v ^{-1}\]

OpenStudy (sirm3d):

hmm, a bernoulli equation with n = 2.

OpenStudy (sirm3d):

the right side is \[-\frac{ e^{2x} }{ x }\]

OpenStudy (sirm3d):

oh well, i'll take it from here. divide by \(xy^2\)\[\huge \frac{ y^{-2}dy }{ dx }-\frac{ 1 }{ x } \underbrace{y^{-1}}_u = -\frac{e^{2x}}{x}\]

OpenStudy (sirm3d):

\[u=y^{-1}\]\[du=-y^{-2}dy\] \[\large -\frac{ du }{ dx }-\frac{ 1 }{ x }u=-\frac{ e^{2x} }{ x }\]

OpenStudy (sirm3d):

\[\frac{ du }{ dx }+\frac{ 1 }{ x }u=\frac{ e^{2x} }{ x }\]\[u(x)=\int\limits x \left( \frac{ e^{2x} }{ x } \right)dx \]

OpenStudy (anonymous):

That'll give \[\frac{ 1 }{ y}=\frac{ 1 }{ 2}e ^{2x}\]

OpenStudy (anonymous):

forgot +c

OpenStudy (anonymous):

then \[y= \frac{ 2 }{ e ^{2x}}+c _{2}\]

OpenStudy (sirm3d):

\[u(x)=\frac{ 1 }{ y }x\], its multiplication, not function notation. apologies.

OpenStudy (sirm3d):

\[\frac{ x }{ y }=\frac{ e^{2x} }{ 2 }+C\]

OpenStudy (anonymous):

yep, thx, was juss rechecking noticed ma mistake

OpenStudy (anonymous):

Thanks a million

OpenStudy (anonymous):

forgot to ask if they solve do i have to give in terms of y or i can leave it in implicit form

OpenStudy (sirm3d):

i would write the final answer free of denominators. \[2x=y e^{2x} + 2Cy\]

OpenStudy (anonymous):

thx

OpenStudy (sirm3d):

yw